# Forecasting Influenza Hospitalizations Using a Bayesian Hierarchical Nonlinear Model with Discrepancy

### v1

@@ 3.8 Parameter estimation and posterior predictive sampling
The models were fit via Markov chain Monte Carlo (MCMC) sampling using the cmdstanr package which was developed and is maintained by the Stan Development Team Stan Development Team (2024) Gabry et al. (2024). Stan implements Hamiltonian Monte Carlo (HMC) sampling with the No-U-turn sampler Hoffman et al. (2014). The cmdstanr package provides several diagnostic statistics for assessing the sampler. As mentioned, most model parameter prior distributions were intended to be uninformative. Plots of posterior distributions for select parameters from ILI and hospitalization models are shown in the supplementary material Wadsworth and Niemi (2024).
We assessed the model fit for four ILI models. These included the SIR and ASG models and models with and without discrepancy modeling. When discrepancy is included, the models are denoted as SIRD and ASGD. These models were fit using US national data from 2010 to 2023 flu seasons, where data from the 2020 season was excluded because of the unique behavior during that season. Assessment for hospitalization modeling was done for six different models. These include NORM, LNORM, and LST models, and models where a quadratic ILI term is included or excluded. To assess posterior sampling convergence, models were fit to data where ILI and hospitalization data up to week 14 of the 2023 season was included. Sampling was done with four chains where from each chain 60,000 posterior draws were sampled, and the first 10,000 draws were discarded as a burn-in. The $\hat{R}$ statistic Vehtari et al. (2021) and the effective sample size ($ESS$) Gelman et al. (2013) were calculated for each parameter. Tables 1 and 2 summarize the maximum $\hat{R}$ and the minimum $ESS$ over all parameters for ILI and hospitalization models respectively. Forecast models for all other weeks of the season were fit using one chain of 60,000 draws where the first 10,000 draws were discarded as a burn-in. For parameters of the ASG models that were prone to cause trouble in posterior sampling, the starting values were set to be the MLEs.
Table 1: Maximum $\hat{R}$ and minimum $ESS$ over all parameters for four ILI models fitted on US data for week 14 of the 2023 season
ASG ASGD SIR SIRD
$\hat{R}$ $<1.001$ $<1.001$ $<1.001$ $1.001$
$ESS$ 72,057 9,980 17,745 6,259
Table 2: Maximum $\hat{R}$ and minimum $ESS$ over all parameters for six hospitalization models fitted on US data for week 14 of the 2023 season
NORM NORM2 LNORM LNORM2 LST LST2
$\hat{R}$ $<1.001$ $<1.001$ $<1.001$ $<1.001$ $<1.001$ $<1.001$
$ESS$ 71,266 73,554 46,747 54,975 7,660 62,663
To obtain forecast distributions of hospitalizations, draws from the posterior predictive distribution from the ILI model were used in conjunction with the posterior distribution of the hospitalizations model. When fitting model (6), MCMC samples of $\widetilde{ILI}_{s,w:(w+4)}$ were saved. Model (7) was fit and MCMC samples for the marginal distributions for the model parameters were saved. To obtain forecast distributions for $H_{s,w+i}$ where $i\in\{1,2,3,4\}$, the following steps are repeated $K$ times where $K$ is an integer for the number of desired samples. We set $K=50,000$.
@@ 5. Step 5:
The sample $\{H^{*}_{s,w+i}\}^{K}$ was then used as the probabilistic forecast for hospitalizations at week $w+i$. For the forecast competition analysis in section 5, all negative values of $\{H^{*}_{s,w+i}\}^{K}$ were set to 0 to reflect realistic values of hospitalizations and comply with the FluSight forecasting rules.
@@ 4 Simulation Study
In this section, we present a simulation study conducted for comparing ILI models and further assessing the hospitalization forecast model. US ILI data is used, and hospitalization data is simulated. A leave-one-season-out (LOSO) approach was combined with a Monte Carlo simulation approach. For each replication, we simulated log-hospitalizations for all weeks during seasons 2010, …, 2022, excluding 2020, using the existing ILI data as a predictive covariate. Each season was in turn “left-out” and treated as if it was the most recent season which we desired to forecast. Fitting and forecasting was then done for weeks 14, 20, 26, 32, and 38 of the left out season, giving two weeks that tend to occur as flu cases increase, two as cases decrease, and one that occurs when the cases may be increasing or decreasing. Week 20 is a week leading up to the holiday week 22 where ILI typically has a local peak. We were particularly interested in how important modeling discrepancy is for forecasting at week 20.
For the simulation of hospitalizations, the parameters $\bm{\alpha}_{s}=(\alpha_{0s},\alpha_{1s},\alpha_{2s})$ and $\sigma^{2}_{\epsilon_{s}}$ from the hospitalization model in (7) were considered the same across all seasons so that all $\bm{\alpha}_{s}=\bm{\alpha}$. The values for $\bm{\alpha}$, $\sigma^{2}_{\epsilon}$, and $\phi$ were estimated by fitting model (7) using ILI and hospitalization data from the 2022 season. For fitting, the hospitalization data was first log-transformed. We took $\bm{\alpha}_{\phi}=(\bm{\alpha},\phi)$ and assigned the noninformative prior $p(\bm{\alpha}_{\phi},\sigma^{2}_{\epsilon})\propto 1/\sigma^{2}_{\epsilon}$. The marginal posterior distribution $\bm{\alpha}_{\phi}|\sigma^{2}_{\epsilon},\bm{H_{22}}$ was then the established posterior multivariate normal distribution and $\sigma^{2}_{\epsilon}|\bm{H_{22}}$ the inverse-$\chi^{2}$ posterior distribution Gelman et al. (2013). The posterior means of those parameters were used as the values from which log-hospitalizations were simulated. The number of replicates in the simulation was 500.
@@
[…]
Figure 7 shows boxplots of the CRPS for the 500 simulation replicates faceted by week and horizon. Notable plots here are for week 20 where the two models accounting for the ILI discrepancy, ASGD and SIRD, are the best performing forecast models, with lower medians and lower variation than the two models not accounting for ILI discrepancy. This makes sense given the forecasts at week 20 are forecasting weeks, 22 and 23 where figure 5 shows a seasonal peak and trough. This confirms that there is value in modeling discrepancy at least during around the holiday weeks near week 22. Among the four models, ASG tends to have the lowest CRPS values though this is not always the case. This suggests that the additional flexibility of the five parameter ASG function is often better for forecasting than that of the three parameter SIR function. Additional plots in the supplemtary material Wadsworth and Niemi (2024) further show that the ASG models often but not always outperform the SIR models according to the CRPS and another common proper scoring rule known as the logarithmic score.
Figure 7: Boxplots of continuous ranked probability score (CRPS) for the four ILI models over all seasons in the simulation study faceted by horizon (x-axis) and week (y-axis). Horizons include 1-4 week ahead forecasts and weeks include weeks 14, 20, 26, 32, and 38 of the flu season
5 Analysis of forecasts for 2023 flu season
In this section we apply the forecast models to make forecasts of the 2023 flu season weekly hospitalizations. The scoring of the forecasts is in the context of the FluSight competition where each competing forecast was submitted as a set of 23 quantiles corresponding to the given probability levels $(0.010,0.025,0.050,0.100,0.150,…,0.950,0.975,\\ 0.990)$. A single forecast is thus comprised of 11 predictive intervals and a median. Forecasts of 1, 2, 3, and 4-week ahead hospitalization counts were requested, and forecasts were made at the state and national level. The first week of forecasting took place during the week of October 7, 2023, and the final week was the week of April 27, 2024 making 29 total weeks of forecasts. The same format was used during the 2021 and 2022 seasons and for the COVID-19 Forecast Hub Mathis et al. (2024); Bracher et al. (2021). Primary scores for evaluating each forecast were the weighted interval score (WIS), the log-weighted interval score (LWIS), and the relative weighted interval score (RWIS). The focus in this section will be on the LWIS (see supplementary material for RWIS based results Wadsworth and Niemi (2024)).
The WIS is a proper scoring rule used for scoring quantile or interval forecasts Gneiting and Raftery (2007); Gneiting and Katzfuss (2014); Bracher et al. (2021) and is defined in (9) where $Q$ is a forecast represented by all included quantiles, $B$ is the number of intervals, $y^{*}$ is the observed value targeted by the forecast, $w_{0}=1/2$ and $w_{b}=\alpha_{b}/2$ are weights for each interval, and $\alpha_{b}$ is the nominal level of the $b^{th}$ interval. $IS_{\alpha}$ is the interval score (IS), a proper scoring rule for a single interval as defined in (10). The goal of the forecaster is to minimize the WIS. The LWIS is the same as the WIS except that it is evaluated over the log of quantiles and the log of the observed value.
@@ 5 Analysis of forecasts for 2023 flu season
[…]
$\beta_{s}$ $\sim N^{+}(0.8,0.3)$
$\rho_{s}$ $\sim N^{+}(0.68,0.08)$
For the ASG model, the MLE $\hat{\theta}_{s}=(\hat{\lambda}_{s},\hat{\eta}_{s},\hat{\mu}_{s},\hat{\sigma}_{1s}^{2},\hat{\sigma}_{2s}^{2})$ was calculated and $\hat{\lambda}_{s}$ was accepted as a fixed value. The remaining estimates were used as starting values for posterior sampling. The transformation $T(\theta_{s})=(\text{log}(\eta_{s}),\mu_{s},\text{log}(\sigma_{1s}^{2}),\text{log}(\sigma_{2s}^{2}))$ was made, and a prior distribution was assigned to $T(\theta_{s})$. The prior distributions for the ILI model under the ASG function are shown in (12). These are slightly informative priors because for most parameters we have an idea what reasonable values may be. Here $m=(0.3,23,3.69,4.7)$ and $C=\text{diag}(0.2,5,2,2)$ where $\text{diag}(\cdot)$ is the diagonal matrix for the given entries.
$T(\theta_{s})$ $\overset{ind}{\sim}MVN(\theta,\Sigma)$ (12)
$T(\theta)$ $\overset{ind}{\sim}MVN(m,C)$
$\Sigma$ $=\text{diag}(\zeta^{2}_{1},...,\zeta^{2}_{4})$
$\zeta_{i}$ $\overset{ind}{\sim}N^{+}(0,4^{2})$
Parameters shared by both the SIR and ASG models are the scale parameter $\kappa_{s}$ and the discrepancy parameters $\sigma_{\gamma}^{2}$ and $\sigma_{\gamma_{W}}^{2}$. These priors are in (13).
$\kappa_{s}$ $\overset{ind}{\sim}N^{+}(0,10,000^{2})$ (13)
$\sigma_{\gamma}^{2}$ $\sim N^{+}(0,.02^{2})$
$\sigma_{\gamma_{W}}^{2}$ $\sim N^{+}(\hat{\sigma}_{W}^{2},1^{2})$
Because of the limited information for estimating $\sigma_{\gamma_{W}}^{2}$, we selected an informative prior distribution by first estimating $\hat{\sigma}_{\gamma_{W}}^{2}$. This was estimated for a given state by first calculating the MLE for $\theta_{s}$. $\widehat{ILI}_{s,w}$ was then predicted such that $\text{logit}(\widehat{ILI}_{s,W})=f_{\hat{\theta}_{s}}(W)$ for each season. Then $\hat{\sigma}_{\gamma_{W}}^{2}$ was calculated as the estimated variance over seasons of the differences $\text{logit}(\widehat{ILI}_{s,w})-\text{logit}(ILI_{s,w})$. When $f_{\theta_{s}}(w)$ was the SIR function, $\widehat{\theta}_{s}$ was calculated using the mle2 function in the bblme package Bolker and R Development Core Team (2023) and the ode function in the deSolve package Karline Soetaert et al. (2010) function in R. Where the ASG function was used, $\hat{\theta}_{s}$ was calculated using the optim function. The prior for the hospitalization models are in (14).
$\alpha_{0s}$ $\overset{ind}{\sim}N(0,5^{2})$ (14)
$\alpha_{1s}$ $\overset{ind}{\sim}N(0,5^{2})$
$\alpha_{2s}$ $\overset{ind}{\sim}N(0,5^{2})$
[…]
$\sigma_{\epsilon_{s}}$ $\overset{ind}{\sim}N^{+}(0,4^{2})$
$\omega_{s}$ $\overset{ind}{\sim}N^{+}(0,15^{2})$
@@
Figure 8: Forecasts 1-4 weeks ahead for US hospitalizations during the 2023 season for weeks 14, 20, 26, and 32. Forecasts are separated by ILI model, and the hospitalization models are all normally distribution. The figure includes hospitalization forecasts where ILI is a linear predictor (left) and where ILI is a quadratic predictor (right). 50% predictive intervals are pink and 95% predictive intervals are red.
Figure 8 shows 1-4-week ahead forecasts for US flu hospitalizations during the 2023 season under the NORM hospitalization model. The forecasts shown are for weeks 14, 20, 26, and 32. The predictive bands are the 50% and 95% predictive intervals. These plots show a tendency to often underpredict hospitalizations. The quadratic models appear to do a better job predicting hospitalizations at the season peak but a poorer job predicting after the peak, whereas the linear models seem to predict well or slightly overpredict after the peak.
Figure 9 shows model performance by LWIS for each week of the season for all 24 models of US hospitalizations. LWIS scores are grouped by ILI model, hospitalization model distribution, and by the linear or quadratic modeling. Here, the smaller the LWIS the better. The weeks around the holiday week 22 are highlighted by a grey band. In several cases, there appears to be a turning point in performance at or near week 22. For most ASG models, the models which include discrepancy tend to outperform those which do not beginning near week 22 and for the rest of the season. For the SIR models, the discrepancy does not always lead to better forecasts, but in for the LNORM hospitalization models, including discrepancy greatly improves forecasts around the holiday week 22. The LNORM LWIS scores tend to be higher than for LST and NORM hospitalization models, but figure 10 and table 3 suggest that the LNORM hospitalization model outperforms LST and NORM for most states and territories.
Figure 9: Each plot shows the log weighted interval score (LWIS) for every week of the 2023 flu season with scores for models including and exluding discrepancy in the ILI model. Scores are separated by hospitalization distribution family and by ILI as a linear or quadratic predictor. Scores for models with an ASG ILI model are above while those with an SIR model are below. The lower the LWIS the better the forecast.
Figure 10: Each plot shows the log weighted interval scores (LWIS) for all 50 US states, PR, DC, and national level forecasts at each week during the 2023 flu season. Scores are averaged over all horizons 1-4 weeks ahead. Scores are faceted by hospitalization model distribution. The lighter the shade, the lower the LWIS with low LWIS being better.
Figure 10 also shows weekly forecast performance by LWIS, but all 53 locations are included. The LWIS is shown across the whole season faceted by hospitalization model distribution. Overall, the weeks leading up to week 22 are the most difficult to forecast, but the LNORM model appears to perform the better than the other models during the weeks around the holiday week 22. The overall, calculated as the mean LWIS over all locations and weeks, for all 24 models is shown in table 3. The first main takeaway is that the top four performing models are SIR ILI models with the LNORM hospitalization models. The next main takeaway is that overall the models which include discrepancy outperformed the models without discrepancy.
Table 3: Overall scores for each of the 24 forecast models. The overall score is the log weighted interval score (LWIS) averaged over all locations, weeks, and horizons. The scores in the first two rows are for linear models, and the scores in the third and fourth rows are for quadratic models. The lowest WISs are bolded.
ASG SIR
LNORM LST NORM LNORM LST NORM
Linear No Disc 0.366 0.392 0.365 0.355 0.416 0.394
Disc 0.358 0.391 0.365 0.343 0.382 0.354
Quadratic No Disc 0.366 0.398 0.377 0.355 0.401 0.377
Disc 0.358 0.390 0.369 0.343 0.378 0.356
It should not be assumed that these two takeaways will apply for future flu seasons. As suggested by the simulation study in the previous section, it may often be the case that the ASG ILI model is better for forecasting. Indeed, a close examination of figures 9 and 10 shows that it is common for the ASG models to show better forecasting skill than the SIR models.

### v2

@@ 3.8 Parameter estimation and posterior predictive sampling
The models were fit via MCMC sampling using the cmdstanr package which was developed and is maintained by the Stan Development Team Stan Development Team (2024) Gabry et al. (2024). Stan implements Hamiltonian Monte Carlo (HMC) sampling with the No-U-turn sampler Hoffman et al. (2014). The cmdstanr package provides several diagnostic statistics for assessing the sampler. Plots of prior and posterior distributions for select parameters from ILI and hospitalization models are included in the supplementary material Wadsworth and Niemi (2024).
Model fit was assessed for four ILI models and for the hospitalization model. The ILI models included the SIR and ASG models and models with and without discrepancy modeling. When discrepancy is included, the models are denoted as SIRD and ASGD. These models were fit using US national data from 2010 to 2023 flu seasons, where data from the 2020 season was excluded because of the unique behavior during that season. Sampling was done with four chains where from each chain 60,000 posterior draws were sampled, and the first 10,000 draws were discarded as a burn-in. The $\hat{R}$ statistic Vehtari et al. (2021) and the effective sample size ($ESS$) Gelman et al. (2013) were calculated for each parameter, and for all models fit, the largest $\hat{R}$ was below the recommend threshold of 1.05, and the smallest $ESS$ was over 500.
To obtain forecast distributions of hospitalizations, draws from the posterior predictive distribution from the ILI model were used in conjunction with the posterior distribution of the hospitalizations model. When fitting model (7), MCMC samples of $\widetilde{ILI}_{s,w:(w+4)}$ were saved. Model (6) was fit and MCMC samples for the marginal distributions for the model parameters were saved. To obtain forecast distributions for $H_{s,w+i}$ where $i\in\{1,2,3,4\}$, the following steps were repeated $K$ times where $K$ is an integer for the number of desired samples. We set $K=50,000$.
@@ 5. Step 5:
The sample $\{H^{*}_{s,w+i}\}^{K}$ was then used as the probabilistic forecast for hospitalizations at week $w+i$. Occasionally $\{H^{*}_{s,w+i}\}^{K}$ included a small number of negative values, which do not make sense when the distribution is meant to forecast hospitalizations, a nonnegative number. We attempted to alleviate this problem by modeling log-hospitalizations or by assuming hospitalizations followed a distribution truncated at 0, but these models led to poorer forecasts and issues drawing from the posterior distribution. For the forecast competition analysis in section 5, all negative values of $\{H^{*}_{s,w+i}\}^{K}$ were set to 0 to reflect realistic values of hospitalizations and comply with the FluSight forecasting rules.
@@ 4 Simulation Study
In this section, we present a simulation study conducted for comparing ILI models and further assessing the hospitalization forecast model. US ILI data is used, and hospitalization data is simulated. A leave-one-season-out (LOSO) approach was combined with a Monte Carlo simulation approach. For each replication, we simulated log-hospitalizations for all weeks during seasons 2010, …, 2022, excluding 2020, using the existing ILI data as a predictive covariate. Each season was in turn “left-out“ and treated as if it was the most recent season for which forecasts are desired. Fitting and forecasting was then done for weeks 14, 20, 26, 32, and 38 of the left out season, giving two weeks that tend to occur as flu cases increase, two as cases decrease, and one that occurs when the cases may be increasing or decreasing. Week 20 is a week leading up to the holiday week 22 where ILI typically has a local peak. We were particularly interested in how important modeling discrepancy is for forecasting at week 20.
Besides the 4 ILI models, we also included forecasts for a naive baseline (BASE) model and an autoregressive integrated moving average (ARIMA) model. The BASE model is a simple random walk with drift model, and the ARIMA model is fit by selecting the best possible model give the data according to a fitting criterion. Both the BASE and ARIMA ILI models were fit using the forecast package in R Hyndman and Khandakar (2008).
For the simulation of hospitalizations, the parameters $\bm{\alpha}_{s}=(\alpha_{0s},\alpha_{1s},\alpha_{2s})$ and $\sigma^{2}_{\epsilon_{s}}$ from the hospitalization model in (6) were considered the same across all seasons so that all $\bm{\alpha}_{s}=\bm{\alpha}$. Note that this is not a model assumption we make but only a simplyfying assumption for simulating the data. The values for $\bm{\alpha}$, $\sigma^{2}_{\epsilon}$, and $\phi$ were estimated by fitting model (6) using ILI and hospitalization data from the 2022 season. For fitting, the hospitalization data was first log-transformed. We took $\bm{\alpha}_{\phi}=(\bm{\alpha},\phi)$ and assigned the noninformative prior $p(\bm{\alpha}_{\phi},\sigma^{2}_{\epsilon})\propto 1/\sigma^{2}_{\epsilon}$. The marginal posterior distribution $\bm{\alpha}_{\phi}|\sigma^{2}_{\epsilon},\bm{H_{22}}$ was then the established posterior multivariate normal distribution and $\sigma^{2}_{\epsilon}|\bm{H_{22}}$ the inverse-$\chi^{2}$ posterior distribution Gelman et al. (2013). The posterior means of those parameters were used as the values from which log-hospitalizations were simulated. The number of replicates in the simulation was 500.
@@
The forecasts used in this section and in the subsequent section follow the FluSight convention of being quantiles corresponding to the 23 given probabilities (0.010,0.025,0.050,0.100,0.150,,0.950,0.975,0.990)\\ (0.010,0.025,0.050,0.100,0.150,…,0.950,0.975,0.990) making each forecast a set of 11 predictive intervals and a predictive median. Figure 7 gives three plots showing results of the study. Plot a) shows WIS scores averaged over each left out season and each of the 500 simulation replicates. The ASGD forecasts on average clearly outperformed the other models considered in almost every week. The ASGD does especially well at week 20 which contains forecasts of the holiday week 22. This demonstrates the particular importance of modeling discrepancy at and around that week. The SIRD outperforms the SIR model, again showing the importance of modeling discrepancy, but it is still outperformed by the ASG and even for a few weeks by the BASE and ARIMA models. Plot b) shows boxplots of all forecast scores by model and separated by horizon forecasts of 1 to 4 weeks ahead. Here also the ASGD performs the best with the other variations of the nonlinear models not always outperforming the BASE or ARIMA models. To assess calibration, plot c) shows the empirical coverage of each model for 11 predictive intervals from 10% to 98% nominal levels. Each model is undercalibrated, but the ASGD model is the closest to being well calibrated, especially at the hightest nominal levels. Table 1 shows summaries of the mean WIS and coverage for 50% and 95% predictive intervals.
Figure 7: Results of forecasts for simulated hospitalizations for ASGD, ASG, SIRD, SIR, BASE, and ARIMA models. Plot a), WIS averaged over all left out seasons and 500 simulation replicated for each of weeks 14, 20, 26, 32, and 38, coloured by model. Plot b), boxplots of all WIS scores separated by forecast horizon of 1 to 4 weeks ahead. Plot c), empirical calibration of each forecast model as the overall forecast coverage of predictive intervals for the 11 nominal levels 10%, 20%, …, 90%, 95%, 98%.
Table 1: Summary values of forecast performance in simulation study for ASGD, ASG, SIRD, SIR, BASE, and ARIMA models. WIS, 50% and 95% predictive interval coverage averaged over all seasons and 500 simulation replicates.
Model WIS 50%Coverage 95% Coverage
ASGD 0.02 0.42 0.91
ASG 0.03 0.33 0.71
ARIMA 0.03 0.35 0.68
BASE 0.03 0.33 0.65
SIRD 0.03 0.33 0.78
SIR 0.05 0.20 0.51
The results of this simulation study demonstrate the effectiveness of using the assymetric Gaussian function along with modeling discrepancy for modeling ILI and making flu hospitalization forecasts in the US. The following section shows similar effectiveness of coupling an effective ILI flu model with a hospitalization model to forecast hospitalizations during the 2023 flu season.
5 Analysis of forecasts for 2023 flu season
In this section we apply the forecast models to make forecasts for the 2023 flu season weekly hospitalizations and compare the results with those of forecasts submitted to FluSight. The scoring of the forecasts is in the context of the FluSight competition where each competing forecast was submitted as a set of 23 quantiles corresponding to given probabilities. Forecasts of 1, 2, 3, and 4-week ahead hospitalization counts were requested, and forecasts were made at the state and national levels. The first week of forecasting took place during the week of October 7, 2023, and the final week was the week of May 4, 2024 making 30 total weeks of forecasts. The same format was used during the 2021 and 2022 seasons and for the COVID-19 Forecast Hub Mathis et al. (2024); Bracher et al. (2021). Primary scores for evaluating each forecast were the WIS and the log-weighted intveral score (LWIS). The LWIS is defined the same as the WIS but where the log of the observation as well as the log forecast quantiles are taken prior to calculating the score. We primarily use the LWIS in this section, though we including results for the WIS in the supplementary materials.
Figure 8 shows hospitalization forecasts at four different weeks during the season for ASGD, ASG, SIRD, and SIR ILI models where hospitalizations are modeled according to model (6) where ILI is a quadratic predictor of hospitalizations, and hospitalizations are assumed to be normally distributed. Of these forecasts, the ASGD appears to do the best job at capturing the trajectory of hospitalizations as well as most frequently making the sharpest predictions, particularly later in the season.
@@ 5 Analysis of forecasts for 2023 flu season
Figure 8: Forecasts 1-4 weeks ahead for US hospitalizations during the 2023 season for weeks 14, 20, 26, and 32. Forecasts are separated by ILI model, and the hospitalization models are all normally distributed and ILI is a quadratic predictor.
For the remainder of this section, focus will be made on forecasts made by coupling the ASGD and SIRD models with the normal hospitalization model where ILI is a quadratic predictor. These forecast schemes are denoted as ASGD_NORM2 and SIRD_NOMR2. In the supplementary materials, we include analysis of 2023 hospitalization forecasts for several other forecast models including ASG and SIR ILI models coupled with other variations of the hospitalization model where the ASGD_NORM2 showed the best results and the SIRD_NORM2 performed nearly as well as the ASGD_NORM2.
The FluSight models which are used for comparison are the 20 non-ensemble models submitted to FluSight during the 2023 season. The submitted forecasts may be found at Github (2024). Among the models was the FluSight baseline. The baseline forecast had as a median the most recently observed hospitalization count, and the forecast uncertainty was based on differences between previous hospitalizations. The baseline forecast model was similar to the baseline forecast model used in previous flu forecasting seasons and in the COVID-19 hub Mathis et al. (2024); Cramer et al. (2022).
Figure 9 shows forecast results for the 2023 flu hospitalization forecasts for all 22 models being compared but with results for the ASGD_NORM2, SIRD_NORM2, and FluSight baseline models highlighted. For each week of the 2023 season, the LWIS averaged over location and horizon for each model is shown in a). Boxplots of LWIS for 1, 2, 3, and 4-week ahead horizons for the ASGD_NORM2, SIRD_NORM2, and baseline models are shown in b). And overall empirical coverage of the 11 predictive intervals for all models is shown in c) with ASGD_NORM2, SIRD_NORM2 and the baseline model highlighted. These plots show that over the season the ASGD_NORM2 model tended to outperform most others in terms of LWIS, and in general is superior to the baseline model. The coverage plot shows that the ASGD_NORM2 model is well calibrated at most levels, with slight overcoverage at the lower levels up to about 80%, and has better coverage than all other models. Similar results for the WIS are shown in the supplementary materials.
Figure 9: a) Log-weighted interval score (LWIS) averaged over location and horizon for each week during the season for the 21 models being compared. Each grey line represents one model, but the ASGD_NORM2 (blue), SIRD_NORM2 (green), and FluSight-baseline (red) are colored to stand out. b) Boxplots of all LWIS scores for 1, 2, 3, and 4-week ahead horizons for the ASGD_NORM2 (blue) and FluSight-baseline (red) models. c) Plot of empirical coverage against theoretical coverage for 11 predictive intervals from the 10% nominal level to the 98%. The black line $x=y$ shows where optimal coverage would fall. Each grey line represents one model, but the ASGD_NORM2 (blue), SIRD_NORM2 and FluSight-baseline (red) are colored to stand out.
During the FluSight competition, most participating teams would occasionally miss a forecast submission for some week, location, or horizon. This makes a simple average of scores an imperfect measure for comparison, as forecasts of certain targets included for one model may be excluded by another model. Thus in figure 10 the ASGD_NORM2 model is compared with all other models individually using LWIS scores averaged over only the forecasts shared by the two models. Each tile represents the ratio of average ASGD_NORM2 LWIS to that of one FluSight model by location or by week. We call this the relative LWIS (RLWIS). A value smaller than 1 indicates the ASGD_NORM2 model outperforms the competing model. A value higher than 1 indicateds the ASGD_NORM2 is outperformed. A grey tile indicates the competing model had no forecasts matching the specified target. Similar plots showing WIS scores are found in the supplementary materials. The ASGD_NORM2 outperforms most other competing models more often than not, but there are some weeks or locations where forecasting was poorer. For example, the plot shows the ASGD_NORM2 performed poorly in Montana relative to most other models.
Table 2 gives summary results for mean and median WIS and LWIS scores, predictive coverage, and relative RWIS and RLWIS scores for each model. Along with the plots in this section, it shows that over the course of the 2023 season, the ASGD_NORM2 outperformed every other model except the UGA_flucast-INFLAenza. The SIRD_NORM2 is also one of the top performing forecast models.
Figure 10: Ratio of average log-weighted interval score (LWIS) for ASGD_NORM2 to that of each competing FluSight model. Scores are averaged over all targets within indicated location (top) or week (bottom) and only include forecasts shared by both models. Lower scores (blue) indicate superior performance by the ASGD_NORM2 model and higher scores (red) indicate superior performance by competing model.
Table 2: Overall summary scores for each of the 20 non-ensemble FluSight models and the ASGD_NORM2 model. Summaries include mean weighted interval and log-weighted interval scores (MWIS/MLWIS), median weighted interval and log-weighted interval scores (MedWIS/MedLWIS), mean squared error difference between predictive model empirical and theoretical coverage (MSEC) see figure 9 c)), percentage of forecasts for all targets during the 2023 season (% Forcs), and relative weighted interval and log-weighted interval scores (RWIS/RLWIS).
MWIS MLWIS MedWIS MedLWIS MSEC % Forcs RWIS RLWIS
ASGD_NORM2 62 0.34 14 0.23 0.001 100 1 1
PSI-PROF 63 0.35 15 0.23 0.003 100 0.90 0.97
SIRD_NORM2 67 0.36 16 0.26 0.001 100 0.93 0.95
CEPH-Rtrend_fluH 69 0.39 13 0.28 0.004 100 0.81 0.86
FluSight-baseline 85 0.41 16 0.32 0.023 100 0.63 0.82
UM-DeepOutbreak 83 0.46 21 0.39 0.010 99 0.64 0.72
cfarenewal-cfaepimlight 76 0.37 15 0.26 0.001 99 0.72 0.92
MOBS-GLEAM_FLUH 66 0.35 15 0.22 0.021 98 0.87 0.95
LosAlamos_NAU-CModel_Flu 242 0.64 24 0.42 0.153 93 0.21 0.53
UGuelph-CompositeCurve 132 0.52 25 0.41 0.018 93 0.43 0.62
LUcompUncertLab-chimera 84 0.43 17 0.33 0.004 92 0.67 0.71
UGA_flucast-INFLAenza 74 0.32 15 0.20 0.003 90 0.77 1.07
cfa-flumech 101 0.41 17 0.27 0.019 89 0.58 0.85
UGA_flucast-Copycat 76 0.35 16 0.21 0.012 87 0.81 0.98
GT-FluFNP 105 0.48 18 0.38 0.046 86 0.45 0.73
UNC_IDD-InfluPaint 107 0.35 16 0.24 0.043 78 0.58 0.95
NU_UCSD-GLEAM_AI_FLUH 98 0.39 22 0.24 0.014 74 0.68 0.83
SigSci-CREG 59 0.49 12 0.31 0.018 71 0.62 0.74
PSI-PROF_beta 77 0.32 18 0.20 0.004 67 0.87 0.93
NIH-Flu_ARIMA 230 0.34 28 0.23 0.001 20 0.57 0.88
GH-model 161 1.52 28 1.47 0.373 17 0.18 0.27
UGA_flucast-OKeeffe 27 0.59 8 0.43 0.003 10 0.60 0.58